OpenStudy (imtant):

How to solve all this three equations to get T1,T2 and P? -T1(4/square root 21)-T2(4/square root 26)+P=0 T1(2/square root 21)-T2(3/square root 26)=0 T1(1/square root 21)+T2(1/square root 26)-981=0..Anyone?

OpenStudy (anonymous):

The roots are there to trick you into thinking it's difficult. For example, add 3*Eq.(3) to Eq.(2) to eliminate T2, then plug T2 into Eq.(2) to find T1 and plug T2 and T1 into Eq.(1) to find P, simple as that!

OpenStudy (imtant):

I don't get it.

OpenStudy (anonymous):

Well, Eq.(3) implies T1*(1/sqrt(21))+T2*(1/sqrt(26))=981, multiply by 3 we'll get T1*(3/sqrt(21))+T2*(3/sqrt(26))=981*3, now when adding the above Eq. to Eq.(2), T2 vanishes, so T1*(5/sqrt(21))=981*3, you can solve for T1 from here, and then use backward substitution to find the others.

OpenStudy (imtant):

Did you mean 'when adding the above Eq. to Eq.(2),' refer to eq 3 and 1 have to added?

OpenStudy (anonymous):

The 'above Eq.' is what you get after multiply the Eq.(3) by 3.

OpenStudy (imtant):

Yes,i get it.Thanks.

OpenStudy (imtant):

So,when i get this 2942.8=-T2(4/squareroot 26)+P,i need to find p first or T2?

OpenStudy (anonymous):

Find T2 first from Eq.(1), because it only involves T1 (which I assume you've ready solved for) and T2, then sub in T1 and T2 into Eq.(1) (as it's the only eq involving P) to find P

OpenStudy (imtant):

Ok,i had already got it.However,thanks,drawar!